A solution containing 0.100 M HF also contains 0.0100 M Ca(NO3)2. Hydrochloric acid is added slowly till the solution has a pH of 2.00. Will a precipitate of calcium fluoride,CaF2 occur? for the HF, Ka=6.46 x 10^-4 and for the CaF2, Ksp=2.69 x10^-11
1
Expert's answer
2019-07-17T06:16:58-0400

CaF2 + 2H3O+ = Ca2+ + 2HF + 2H2O

[H3O+] = 10-pH = 10-2

aF- = Ka, HF / ([H3O+] + Ka, HF) = 6.46 х 10-4 / (10-2 + 6.46 х 10-4) = 6.1 х 10-2

[F-] = C(HF) x aF- = 0.100 x 6.1 х 10-2 = 6.1 х 10-3 M

[Ca2+] x [F-] = 0.0100 x 6.1 х 10-3 = 6.1 х 10-5 > Ksp = 2.69 х 10-11